1. Calculate the theoretical yield for the following reaction. Use 1 g of acetone and an excess of NaBH4. NaBH OH CH3CH2OH
Answer #1

Since excess NaBH4 is used, so, all ketone must react theoretically. Now, from the equation it is clear that, from 1 mol of ketone, 1 mol of corresponding alcohol is produced.

Now, molar mass of acetone (C3H6O) = (3 x 12 + 6 x 1 + 1 x 16) g.mol-1 = 58 g.mol-1

molar mass of isopropanol(C3H8O) = (3 x 12 + 8 x 1 + 1 x 16) g.mol-1 = 60 g.mol-1

So, from 58 g of acetone if 100% yield (theoretical yield) occurs then we should ge 60 g of alcohol. So, from 1 g of acetone we should get (60 / 58) g alcohol = 1.034 g.

So, theoretical yield should be 1.034 g.